OpenStudy (anonymous):

Please help me with this one! Enter here ____________ an algebraic expression that gives the area of a circle with a circumference. make up one that would work. Hint: The circumference of a circle with (r) radius is 2pi(r) , and its area is pi(r)squared

OpenStudy (anonymous):

What do you think your approach to this will be?

OpenStudy (anonymous):

The circumference is C=2pi*r. Area is pi r^2. Thus r=C/2pi and A=pi(C/2pi)^2 = C^2/4pi

OpenStudy (anonymous):

Yes, I'm glad you know how to do this. I was hoping to see if lancer did.

OpenStudy (anonymous):

ok plz give me a medal for my repsond

OpenStudy (anonymous):

polpak

OpenStudy (anonymous):

popak how do you scan something to the computer

OpenStudy (anonymous):

With a scanner. Though you could also just take a picture.

OpenStudy (anonymous):

how i need help

OpenStudy (anonymous):

It's not something I can really help you with.

OpenStudy (anonymous):

with my hw

OpenStudy (anonymous):

i want to attach it hwo

OpenStudy (anonymous):

There are two circles , x^2 + y^2 = 0 and x^2 + y^2 = 4

OpenStudy (anonymous):

is it r8 guys